# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ 2.1.2. Coupling between $\mathrm{SLE}_{4}$ and the GFF
In simply connected domains, there is already a deep connection between the GFF and $\mathrm{SLE}_{4}$ [SS13, Dub09, WP21].
Theorem 7 ($\mathrm{SLE}_{4}$ is a local set for the GFF).
Let $D$ be a simply connected domain with $x\neq y\in\partial D$. There exists a coupling $(\Phi_{D}^{g_{D}},\eta,\Phi_{\eta})$ with $\eta\sim\nu_{D,4}^{\#}(x,y)$ an $\mathrm{SLE}_{4}$ curve from $x$ to $y$ in $D$ that is a local set coupling, with $\Phi_{\eta}$ given by $h_{\eta}$ on $D\setminus\eta$, and $h_{\eta}(z):=-\lambda+2\lambda\mathbf{1}(z\in(D\setminus\eta)_{\text{right}})$.
@@
Proposition 8 (General level-lines, [ASW19, Lemma 15]).
Let $D$ be a non-crossing admissible domain and $g$ a bounded harmonic function in $D$ with Dirichlet boundary conditions $-\lambda$ (resp. $+\lambda$) on $(yx)$ (resp. $(xy)$), arbitrary on other parts of $\partial D$. Then there exists a unique (up to time-reparametrization) law on random simple curves $(\eta(t),0\leq t\leq\tau_{D})$ in $D$ with $\eta(0)=x$ that can be coupled with a GFF $\Phi_{D}^{g}$ such that $(\Phi_{D}^{g},\eta,\Phi_{\eta})$ is a local set coupling, and $\Phi_{\eta}$ corresponds to the unique harmonic function in $D\setminus\eta$ with boundary values given by $g$ on $\partial D$ and $-\lambda$ (resp. $+\lambda$) on the left side of $\eta_{\tau_{D}}$ (resp. on the right side of $\eta_{\tau_{D}}$). We call $\eta$ the generalized level-line of $\Phi_{D}^{g}$.
It is also worth mentioning that the probability that $\tau_{D}<+\infty$ and that $\eta(t)$ converges to a point in $\partial D\setminus\left\{y\right\}$ as $t\to\tau_{D}^{-}$ is equal to 00 if and only if |$\absolutevalue{g}\geq\lambda$ on $\partial D$ [PW17, ASW19]: in this case, the level-line can be parametrized for instance by its capacity in $\mathbb{D}$ seen from $y$, so that $\eta(t)\xrightarrow[t\to\infty]{}y$ a.s. Furthermore, in this coupling $\eta$ is measurable with respect to $\Phi_{D}^{g}$.
In the case of $D$ a crossing domain, generalized level-lines can also be defined for GFFs $\Phi_{D}^{g}$, where this time $g$ is multivalued with boundary values locally constant with a change of ±$\pm 2\lambda$ around $x$ and $y$ on $\partial D$ (see Section 2.3). This was proven for instance in [IK13] for the case of conformal annuli, and extends to the general case by absolute continuity. For the sake of completeness, we include a more thorough explanation in Appendix A.
@@ 2.2. Constructing the non-crossing $\mathrm{SLE}_{4}$ measure
In this section, we construct a measure on curves joining two points in the same boundary component of an $n$-connected domain $D$ (non-crossing case). We assume that no boundary component is reduced to a point. Such domains are conformally equivalent to a non-crossing admissible domain as defined in Section 1.1: we can actually assume that the boundary components different to $\partial\mathbb{D}$ are circles, in particular they are smooth. Because the measure we construct is conformally invariant, we restrict the study to this kind of domain.
For $\mathbf{b}=(b_{1},\dots,b_{n})\in\left\{\pm\lambda\right\}^{n}$, we define the event:
$\mathcal{A}_{\mathbf{b}}:=\left\{\eta\text{ curve from }x\text{ to }y\text{ in}D\,|\,\forall i:b_{i}=-\lambda,\mathcal{B}_{i}\subset(D\setminus\eta)_{\text{left}}\text{ and }\forall i:b_{i}=+\lambda,\mathcal{B}_{i}\subset(D\setminus\eta)_{\text{right}}\right\},$
i.e. the set of curves from $x$ to $y$ in $D$ leaving the components of $\mathcal{B}$ corresponding to the positive components of $\mathbf{b}$ to the left, the others to the right. Note that these events partition $\left\{\eta\subset D\right\}$, since $\mathcal{A}_{\mathbf{b}}\cap\mathcal{A}_{\mathbf{b}^{\prime}}=\emptyset$ for $\mathbf{b}^{\prime}\neq\mathbf{b}$, and that any simple curve from $x$ to $y$ in $D$ must belong to $\mathcal{A}_{\mathbf{b}^{\prime}}$ for some $\mathbf{b}^{\prime}\in\left\{\pm\lambda\right\}^{n}$.
Let also $\mathcal{B}_{\mathbf{b}}^{-}$ (resp. $\mathcal{B}_{\mathbf{b}}^{+}$) correspond to the union of $\mathcal{B}_{l}$ (resp. $\mathcal{B}_{r}$) and of the boundary components $\mathcal{B}_{i}$ such that $b_{i}=-\lambda$ (resp. $b_{i}=+\lambda$).
Figure 1. An example of a domain with the definitions above, together with a test domain $D^{\prime}$ shaded in light blue. Here the curve $\eta$ depicted in the middle is in $\mathcal{A}_{\mathbf{b}}$ for $\mathbf{b}=\left\{-\lambda,+\lambda,+\lambda\right\}$. In particular, the boundary values of $g_{\mathbf{b}}$ are $-\lambda$ and $+\lambda$ exactly on the parts of $\partial D$ that lie the left and on the right of $\eta$, respectively.
Figure 2. An example of two non path-homotopic curves in the event $\mathcal{A}_{\left\{1\right\}}$. In general, there are infinitely (countably) many homotopy classes of distinct curves corresponding to every event $\mathcal{A}_{\mathbf{b}}$.
@@
We define $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ as the law of the generalized level line of a GFF $\Phi_{D}$ in $D$ with boundary conditions given by $g_{\mathbf{b}}$, the harmonic function with boundary values $-\lambda$ on $\mathcal{B}_{\mathbf{b}}^{-}$, $+\lambda$ on $\mathcal{B}_{\mathbf{b}}^{+}$. The important picture to keep in mind is that on the event $\mathcal{A}_{\mathbf{b}}$ and conditionally on $\eta$ the level-line of the GFF, the latter is equal in law to a GFF in $D\setminus\eta$ with piecewise constant mean: $-\lambda$ on the left of $\eta$, $+\lambda$ on its right. That would not be true if $\eta\notin\mathcal{A}_{\mathbf{b}}$.
Theorem 9.
Let $D\subset\mathbb{D}$ be admissible. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(9) $\nu_{D}(\eta):=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}(\eta)\cdot Z_{D,\mathbf{b}},$
@@
(10) $Z_{D,\mathbf{b}}=\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2}),$
is the measure of $\mathrm{SLE}_{4}$ in $(D,x,y)$.
Remark 2.1.
There are several equivalent ways of expressing the mass term $Z_{D,\mathbf{b}}$ (see also Section 2.2). It will appear in the proof that for $D^{\prime\prime}\subset D$ a (non-crossing) test domain, one has
(11) $\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime\prime}}}$ $=\exp(\frac{1}{2}(\norm{g_{D^{\prime\prime}}}_{\nabla(D^{\prime\prime})}^{2}-\norm{g_{\mathbf{b}}}_{\nabla(D)}^{2}))$
well-defined without the need for regularization. Here again this might be rewritten
$\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime\prime}}}=\exp(2\lambda^{2}(M(\mathcal{B}_{l}^{\prime\prime}\cup\mathcal{B}_{l},\mathcal{B}_{r}^{\prime\prime}\cup\mathcal{B}_{r})-M(\mathcal{B}_{\mathbf{b}}^{-},\mathcal{B}_{\mathbf{b}}^{+}))),$
in terms of inverse extremal distance, see Section 1.3.
This is consistent with Equation (10), given that $Z_{D^{\prime\prime}}=H_{D^{\prime\prime}}(x,y)^{h}$ and that in [Dub09] it is argued (proof of Proposition 5.2, page 29) that
$\exp(-\frac{1}{2}\norm{g_{D^{\prime\prime}}}_{\nabla(D^{\prime\prime}),\text{reg}}^{2})=\mathrm{const}\cdot H_{D^{\prime\prime}}(x,y)^{\nicefrac{{1}}{{4}}},$
where $\nicefrac{{1}}{{4}}$ is the boundary scaling exponent $h_{\kappa}$ for $\kappa=4$.
Before going to the proof of the Theorem, we make some simple observations about a GFF in $D$ and its coupling with a level-line. We fix $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$. Let $(\Phi_{D},\eta)$ be a local set with $\Phi_{D}:=\Phi_{D}^{g_{\mathbf{b}}}$ a GFF in $D$ with boundary values given by $g_{\mathbf{b}}$. On the event $\mathcal{A}_{\mathbf{b}}$, the harmonic part $h_{\eta}$ is two-valued and reads
(12) $h_{\eta}(z)=(-\lambda+2\lambda\mathbf{1}_{\text{ right of }\eta}(z)).$
@@
where $\Phi_{D\setminus\eta}^{0}$ should be understood as $\Phi_{(D\setminus\eta)_{\text{left}}}^{0}+\Phi_{(D\setminus\eta)_{\text{right}}}^{0}$. Let us now consider $D^{\prime}\subset D$ a test domain with $\left\{\eta\subset D^{\prime}\right\}\subset\mathcal{A}_{\mathbf{b}}$. Using the standard Markov property on $\Phi_{D\setminus\eta}^{0}$ on the event that $\eta\subset D^{\prime}$, we can further expand:
@@
where $u_{\mathcal{B}^{\prime}}^{D^{\prime}\setminus\eta}$ is a harmonic function in $D^{\prime}\setminus\eta$ corresponding to the harmonic extension of the restriction $\xi_{l}$ (resp. $\xi_{r}$) of $\Phi_{D\setminus\eta}^{0}$ on $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$), and of zero on the other parts of the boundary (including $\eta$). In other words, since $\xi=(\xi_{l},\xi_{r})$ is distributed like a centered Gaussian process on $\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}$ whose covariance matrix is diagonal with elements $(G_{D\setminus\eta})|_{\mathcal{B}_{l}^{\prime}}$ and $(G_{D\setminus\eta})|_{\mathcal{B}_{r}^{\prime}}$, conditionally on $\eta$ still and on the event $\left\{\eta\subset D^{\prime}\right\}\subset\mathcal{A}_{\mathbf{b}}$, we can write $\Phi_{D}|_{D^{\prime}}$ as $\Phi_{D^{\prime}\setminus\eta}^{0}+\bar{h}_{\eta}$, with $\bar{h}_{\eta}$ is a random harmonic function in $D^{\prime}\setminus\eta$ with boundary values
$\bar{h}_{\eta}=\begin{dcases}-\lambda&:\text{ on the left-side of }\eta\\ +\lambda&:\text{ on the right-side of }\eta\\ \xi_{l}\sim\mathcal{N}(-\lambda,G_{D\setminus\eta}|_{\mathcal{B}_{l}^{\prime}})&:\text{ on }\mathcal{B}_{l}^{\prime}\\ \xi_{r}\sim\mathcal{N}(+\lambda,G_{D\setminus\eta}|_{\mathcal{B}_{r}^{\prime}})&:\text{ on }\mathcal{B}_{r}^{\prime}\end{dcases}.$
We consider, similarly, another local set $(\tilde{\Phi}_{D^{\prime}},\tilde{\eta})$ in $D^{\prime}$ with $\tilde{\Phi}_{D^{\prime}}=\tilde{\Phi}_{D^{\prime}}^{g_{D^{\prime}}}$, so that conditionally on $\tilde{\eta}$ we can write $\tilde{\Phi}_{D^{\prime}}=\tilde{\Phi}_{D^{\prime}\setminus\tilde{\eta}}^{0}+h_{\tilde{\eta}}$, with
$h_{\tilde{\eta}}(z)=(-\lambda+2\lambda\mathbf{1}_{\text{ right of }\tilde{\eta}}(z)).$
@@
For computations, it will sometimes be easier to work with a GFF with zero-boundary conditions, coupled with a level-line such that the resulting harmonic part $h_{\tilde{\eta}}$ is no longer only a two-valued function, but is subtracted $g_{D^{\prime}}$. The marginal law of $\tilde{\eta}$ in this coupling will be denoted $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}$ (the dependency in $x$,$y$ is implicit), and in this case is equal to the $\mathrm{SLE}_{4}$ measure $\nu_{D}^{\#}(z,w)$ by Theorem 7 (from now on we will only be concerned with $\mathrm{SLE}_{4}$ and therefore let $\nu_{D}$ implicitly refer to this value of $\kappa$). On the other hand, the marginal law of $\eta$ in the coupling $(\eta,\Phi_{D})$ as presented above on the event $\left\{\eta\subset D^{\prime}\right\}$ and given the values of $\xi$ is denoted $\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$. It therefore corresponds to the law of the level-line of a GFF in $D^{\prime}$ whose mean is the harmonic extension of $\xi$ on $\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}$ and $-\lambda$ (resp. $+\lambda$) on $\mathcal{B}_{l}\setminus\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}\setminus\mathcal{B}_{r}^{\prime}$).
We now turn to the proof of the Theorem, that goes in three steps: we first establish a purely measure-theoretic Lemma, which then allows us to compute the Radon-Nikodym derivative by fixing $\xi$, and then averaging. The two latter steps are done in a Proposition and a Lemma.
Lemma 10.
@@
Let $X,Y_{1},Y_{2}$ be random variables on the same probability space, taking values in a Polish space $E$, with associated distributions $\mathbb{P}_{X},\mathbb{P}_{Y_{1}},\mathbb{P}_{Y_{2}}$ respectively. Denote also by $\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$ the conditional law of $Y_{1}$ on ${Y_{2}=y_{2}}$ (similarly for $\mathbb{P}_{Y_{2}|Y_{1}=x}$) and assume that $\mathbb{P}_{X}\ll\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$, for $\mathbb{P}_{Y_{2}}$-almost every $y_{2}\in E$. The following equality between Radon-Nikodym derivatives holds $\mathbb{P}_{X}$-a.s.
@@
See Appendix B. ∎
Proof of Theorem 9.
We proceed by induction on $n$, the number of connected components of $\mathcal{B}$ that do not contain $x$ and $y$. For clarity we temporarily denote the measures in the statement of the theorem as $\tilde{\nu}_{D}$, as we are yet to prove that they correspond to the $\mathrm{SLE}_{4}$ measures.
When $n=0$, Equation (9) defines $\tilde{\nu}_{D}$ as the law of the level line of a GFF in $D$ with boundary conditions $-\lambda$ on the counterclockwise arc $y\to x$, $+\lambda$ on the other, which corresponds to $\mathrm{SLE}_{4}$ in $D$ upon setting $Z_{D}=H_{D}(x,y)^{\nicefrac{{1}}{{4}}}$.
Let now $n\geq 1$ and $D^{\prime\prime}\subset D$ be a test domain: because $D^{\prime\prime}$ is simply connected by assumption, there exists $\mathbf{b}^{\prime}=(b_{1}^{\prime},\dots,b_{n}^{\prime})\in\left\{\pm\lambda\right\}^{n}$ such that $\left\{\eta\subset D^{\prime\prime}\right\}\subset\mathcal{A}_{\mathbf{b}^{\prime}}$. Up to relabeling the components of $\mathcal{B}$ we may assume that $\mathbf{b}^{\prime}=\left\{-\lambda,\dots,-\lambda,+\lambda,\dots,+\lambda\right\}$ with $1\leq k\leq n$ such that $b_{k}^{\prime}=-\lambda,b_{k+1}^{\prime}=+\lambda$, so that if $D^{\prime}$ is the union of $D^{\prime\prime}$ and the component of $D\setminus D^{\prime\prime}$ on the left of $\eta$, we have by induction that
@@
Figure 3.

### v2

@@ 2.1.2. Coupling between $\mathrm{SLE}_{4}$ and the GFF
In simply connected domains, the coupling between GFF and $\mathrm{SLE}_{4}$ was discovered in [35, 12], see also [43, Chapter 5]. To state this result, recall that we denote by $g_{D}$ the harmonic function with boundary values $\lambda$ (resp. $\lambda$) on $(yx)$ (resp. on $(xy)$), and by $\Phi_{D}^{g_{D}}=\Phi_{D}^{0}+g_{D}$ a GFF in $D$ with mean given by $g_{D}$.
Theorem 2.4 ($\mathrm{SLE}_{4}$ is a local set for the GFF).
Let $D$ be a simply connected domain with distinct $x,y\in\partial D$. There exists a coupling $(\Phi_{D}^{g_{D}},\eta,\Phi_{\eta})$ with $\eta\sim\nu_{D,4}^{\#}(x,y)$ an $\mathrm{SLE}_{4}$ curve from $x$ to $y$ in $D$ that is a local set coupling, with $\Phi_{\eta}$ given by $h_{\eta}$ on $D\setminus\eta$, $h_{\eta}(z):=-\lambda+2\lambda\mathbf{1}(z\in(D\setminus\eta)_{\text{right}})$. Furthermore, in this coupling, $\eta$ is a deterministic function of $\Phi_{D}^{g_{D}}$.
@@
Proposition 2.5 (General level-lines, [4, Lemma 15]).
Let $(D,x,y)$ be a non-crossing admissible domain and $g$ a bounded harmonic function in $D$ with Dirichlet boundary conditions $\lambda$ (resp. $\lambda$) on $(yx)$ (resp. $(xy)$), arbitrary on other parts of $\partial D$. Then there exists a unique law on random simple curves $\eta$ in $D$ (modulo time-reparametrisation, up to its exit time $t_{\eta}$ of $D$) with $\eta(0)=x$ that can be coupled with a GFF $\Phi_{D}^{g}$ such that $(\Phi_{D}^{g},\eta,\Phi_{\eta})$ is a local set coupling, and $\Phi_{\eta}$ corresponds to the unique harmonic function in $D\setminus\eta$ with boundary values given by $g$ on $\partial D$ and $\lambda$ (resp. $\lambda$) on the left side of $\eta$ (resp. on the right side of $\eta$). In this coupling $\eta$ is a deterministic function of $\Phi_{D}^{g}$ and we call it the generalised level-line of $\Phi_{D}^{g}$.
[…]
In the case of $(D,x,y)$ a crossing domain, generalised level-lines can also be defined for GFFs $\Phi_{D}^{g}$, where this time $g$ is harmonic multivalued with boundary values locally constant with a change of $\pm 2\lambda$ around $x$ and $y$ on $\partial D$ (see Section 2.3). The proof is very similar to [4, Lemma 16], but for the sake of completeness we include a more thorough explanation in Appendix A.
@@ 2.2. Constructing the non-crossing $\mathrm{SLE}_{4}$ measure
In this section, we construct a measure on curves joining two points in the same boundary component of a non-crossing admissible domain $(D,x,y)$ as defined in Section 1.1. We start by explaining the set-up, then state and prove Theorem 2.6 and finish the section with some additional remarks.
2.2.1. Set-up and statement
For $\mathbf{b}=(b_{1},\dots,b_{n})\in\left\{\pm\lambda\right\}^{n}$, we consider the set of curves from $x$ to $y$ in $D$ leaving the components of $\mathcal{B}$ corresponding to the negative components of $\mathbf{b}$ to the left, the others to the right. Formally, we define the events
$\mathcal{A}_{\mathbf{b}}:=\{\eta\text{ curve }x\to y\text{ in }D\,|\,\forall j$ $\text{ s.t. }b_{j}=-\lambda,\mathcal{B}_{j}\subset(D\setminus\eta)_{\text{left}}\text{ and }\forall j\text{ s.t. }b_{j}=+\lambda,\mathcal{B}_{j}\subset(D\setminus\eta)_{\text{right}}\},$
with curves here understood modulo time-increasing reparametrisation. These events partition $\left\{\eta\subset D\right\}$ since $\mathcal{A}_{\mathbf{b}}\cap\mathcal{A}_{\mathbf{b}^{\prime}}=\emptyset$ for $\mathbf{b}^{\prime}\neq\mathbf{b}$, and any simple curve from $x$ to $y$ in $D$ must belong to $\mathcal{A}_{\mathbf{b}^{\prime}}$ for some $\mathbf{b}^{\prime}\in\left\{\pm\lambda\right\}^{n}$.
Let also $\mathcal{B}_{\mathbf{b}}^{-}$ (resp. $\mathcal{B}_{\mathbf{b}}^{+}$) correspond to the union of $\mathcal{B}_{l}$ (resp. $\mathcal{B}_{r}$) and of the boundary components $\mathcal{B}_{j}$ such that $b_{j}=-\lambda$ (resp. $b_{j}=+\lambda$).
Figure 1. An example of a domain with the definitions above, together with a test domain $D^{\prime}$ shaded in light blue. Here the curve $\eta$ depicted in the middle is in $\mathcal{A}_{\mathbf{b}}$ for $\mathbf{b}=\left\{-\lambda,+\lambda,+\lambda\right\}$. In particular, the boundary values of $g_{\mathbf{b}}$ are $\lambda$ and $\lambda$ exactly on the parts of $\partial D$ that lie on the left and on the right of $\eta$, respectively.
Figure 2. An example of two non path-homotopic curves in the event $\mathcal{A}_{\left\{-\lambda,+\lambda\right\}}$. In general, there are infinitely (countably) many homotopy classes of distinct curves corresponding to every event $\mathcal{A}_{\mathbf{b}}$.
@@ 2.2.1. Set-up and statement
We define $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ as the law of the generalised level line of a GFF $\Phi_{D}^{g_{\mathbf{b}}}$ in $D$ with boundary conditions given by $g_{\mathbf{b}}$, the harmonic function with boundary values $\lambda$ on $\mathcal{B}_{\mathbf{b}}^{-}$, $\lambda$ on $\mathcal{B}_{\mathbf{b}}^{+}$ (the dependency in $x,y$ is implicit). The important picture to keep in mind is that on the event $\mathcal{A}_{\mathbf{b}}$ and conditionally on $\eta$ the level-line of the GFF, the latter is equal in law to a GFF in $D\setminus\eta$ with piecewise constant mean: $\lambda$ on the left of $\eta$, $\lambda$ on its right. That would not be true if $\eta\notin\mathcal{A}_{\mathbf{b}}$.
Theorem 2.6.
Let $(D,x,y)$ be a non-crossing admissible domain. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(13) $\nu_{D}:=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}\cdot Z_{D,\mathbf{b}},$
@@
(14) $Z_{D,\mathbf{b}}=\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})$
is the conformal restriction measure of $\mathrm{SLE}_{4}$ in $(D,x,y)$.
An immediate corollary is the finiteness of the partition function, with an explicit expression.
Corollary 2.7.
The partition function of $\mathrm{SLE}_{4}$ in a non-crossing admissible domain $(D,x,y)$ can be expressed as
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})\cdot\mathbb{P}_{D}^{g_{\mathbf{b}}}(\mathcal{A}_{\mathbf{b}})$
and it is finite.
Alternatively, Equation (8) can be used to write
$Z_{D}(x,y)=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(\frac{1}{2}(\norm{g_{D^{\prime\prime}}}_{\nabla(D^{\prime\prime}),\mathrm{reg}}^{2}-\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2}))H_{D^{\prime\prime}}(x,y)^{\nicefrac{{1}}{{4}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}(\mathcal{A}_{\mathbf{b}})$
from which conformal covariance is perhaps best read, given that the the probability term and the difference of Dirichlet energies is conformally invariant (see Section 1.3) and that the boundary Poisson kernel is conformally covariant as in Equation (4).
2.2.2. Proof of the theorem
Before going to the proof of Theorem 2.6, we make some simple observations about a GFF in $D$ and its coupling with a level-line. We fix $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$. Let $(\Phi_{D},\eta)$ be a local set coupling with $\Phi_{D}:=\Phi_{D}^{g_{\mathbf{b}}}$ a GFF in $D$ with boundary values given by $g_{\mathbf{b}}$. On the event $\mathcal{A}_{\mathbf{b}}$, the harmonic part $h_{\eta}$ is two-valued and reads
(15) $h_{\eta}(z)=(-\lambda+2\lambda\mathbf{1}_{(D\setminus\eta)_{\text{right}}}(z)).$
@@ 2.2.2. Proof of the theorem
where $\Phi_{D\setminus\eta}^{0}$ should be understood as $\Phi_{(D\setminus\eta)_{\text{left}}}^{0}+\Phi_{(D\setminus\eta)_{\text{right}}}^{0}$. Let us now consider $D^{\prime}\subset D$ a test domain with $\left\{\eta\subset D^{\prime}\right\}\subset\mathcal{A}_{\mathbf{b}}$ (see again Fig 1). Using the standard Markov property on $\Phi_{D\setminus\eta}^{0}$ on the event that $\eta\subset D^{\prime}$, we can further expand:
@@
where $u_{\mathcal{B}^{\prime}}^{D^{\prime}\setminus\eta}$ is a harmonic function in $D^{\prime}\setminus\eta$ corresponding to the harmonic extension of the restriction $\xi_{l}$ (resp. $\xi_{r}$) of $\Phi_{D\setminus\eta}^{0}$ on $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$), and of zero on the other parts of the boundary (including $\eta$). In other words, since $\xi=(\xi_{l},\xi_{r})$ is distributed like a centred Gaussian process on $\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}$ whose covariance matrix is diagonal with elements $(G_{D\setminus\eta})|_{\mathcal{B}_{l}^{\prime}\times\mathcal{B}_{l}^{\prime}}$ and $(G_{D\setminus\eta})|_{\mathcal{B}_{r}^{\prime}\times\mathcal{B}_{r}^{\prime}}$, conditionally on $\eta$ still and on the event $\left\{\eta\subset D^{\prime}\right\}\subset\mathcal{A}_{\mathbf{b}}$, we can alternatively write $\Phi_{D}|_{D^{\prime}}$ as $\Phi_{D^{\prime}\setminus\eta}^{0}+\bar{h}_{\eta}$, with $\bar{h}_{\eta}$ a random harmonic function in $D^{\prime}\setminus\eta$ with boundary values
[…]
$h_{\tilde{\eta}}(z)=(-\lambda+2\lambda\mathbf{1}_{(D^{\prime}\setminus\tilde{\eta})_{\text{right}}}(z)).$
@@
For computations, it will sometimes be easier to work with a GFF with zero-boundary conditions, coupled with a level-line such that the resulting harmonic part $h_{\tilde{\eta}}$ is no longer only a two-valued function, but includes a subtraction by $g_{D^{\prime}}$. The marginal law of $\tilde{\eta}$ in this coupling will be denoted $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}$, and in this case is equal to the $\mathrm{SLE}_{4}$ measure $\nu_{D}^{\#}(x,y)$ by Theorem 2.4.
On the other hand, the marginal law of $\eta$ in the coupling $(\eta,\Phi_{D})$ as presented above on the event $\left\{\eta\subset D^{\prime}\right\}$ and given the values of $\xi$ is denoted $\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$. It therefore corresponds to the law of the level-line of a GFF in $D^{\prime}$ whose mean is the harmonic extension of $\xi$ on $\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}$ and $\lambda$ (resp. $\lambda$) on $\mathcal{B}_{l}\setminus\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}\setminus\mathcal{B}_{r}^{\prime}$).
We now turn to the proof of Theorem 2.6, which essentially consists of leveraging the properties of the GFF to compute the Radon-Nikodym derivatives for the laws of its level-lines. The key computation is contained in Proposition 2.9. We start with a simple measure-theoretic lemma.
Lemma 2.8.
@@
Let $X,Y_{1},Y_{2}$ be random variables on the same probability space, taking values in a Polish space $E$, with associated distributions $\mathbb{P}_{X},\mathbb{P}_{Y_{1}},\mathbb{P}_{Y_{2}}$ respectively. Denote also by $\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$ a version of the conditional law of $Y_{1}$ on $Y_{2}=y_{2}$} (similarly for $\mathbb{P}_{Y_{2}|Y_{1}=x}$) and assume that $\mathbb{P}_{X}\ll\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$, for $\mathbb{P}_{Y_{2}}$-almost every $y_{2}\in E$. Then the following equality between Radon-Nikodym derivatives holds $\mathbb{P}_{X}$-a.s.
@@
See Appendix C. ∎
Proof of Theorem 2.6.
We proceed by induction on $n$, the number of connected components of $\mathcal{B}$ that do not contain $x$ and $y$  [^4] One might wonder why we use induction and do not do the computations directly. The reason is to avoid dealing with two boundaries at once as it makes some of the computations presented below more lengthy and involved.. For clarity we temporarily denote the measures in the statement of the theorem as $\tilde{\nu}_{D}$, as we have not yet proved that they correspond to the $\mathrm{SLE}_{4}$ measures.
When $n=0$, Equation (13) defines $\tilde{\nu}_{D}$ as the law of the level line of a GFF in $D$ with boundary conditions $\lambda$ on the counterclockwise arc $y\to x$, $\lambda$ on the other, which corresponds to $\mathrm{SLE}_{4}$. Furthermore, the expression of $Z_{D}$ in Equation (14) agrees with that of the partition function of $\mathrm{SLE}_{4}$ in simply connected domains since from Equation (8) it follows that
(16) $\exp(-\frac{1}{2}\norm{g_{D}}_{\nabla(D),\mathrm{reg}}^{2})=H_{D}(x,y)^{\nicefrac{{1}}{{4}}}.$
Let now $n\geq 1$ and $D^{\prime\prime}\subset D$ be a test domain: because $D^{\prime\prime}$ is simply connected by assumption, there exists $\mathbf{b}^{\prime}=(b_{1}^{\prime},\dots,b_{n}^{\prime})\in\left\{\pm\lambda\right\}^{n}$ such that $\left\{\eta\subset D^{\prime\prime}\right\}\subset\mathcal{A}_{\mathbf{b}^{\prime}}$. Up to relabeling the components of $\mathcal{B}$ we may assume that $\mathbf{b}^{\prime}=\left\{-\lambda,\dots,-\lambda,+\lambda,\dots,+\lambda\right\}$ with $1\leq k\leq n$ such that $b_{k}^{\prime}=-\lambda,b_{k+1}^{\prime}=+\lambda$ (assuming w.l.o.g. that $k\geq 1$, otherwise reasoning on the right of $D^{\prime\prime}$ instead), so that if $D^{\prime}$ is the union of $D^{\prime\prime}$ and the component of $D\setminus D^{\prime\prime}$ on the left of $\eta$, we have by induction that
@@
Figure 3. $D^{\prime\prime}$ corresponds to the darkest shade, $D^{\prime}$ to the two darker shades (left of orange curve $\mathcal{B}_{r}^{\prime}$) and $D$ to the three shades.